LIFETIMES OF THE Ã STATES OF C_{3}, C_{3}-NE, AND C_{3}-AR

YI-JEN WANG, CHIAO-WEI CHEN, LIUZHU ZHOU, ANTHONY J. MERER, \underline{YEN-CHU HSU}, Institute of Atomic and Molecular Sciences, Academia Sinica, P. O. Box 23-166, Taipei 10617, Taiwan, R. O. C..

The fluorescence lifetimes of the Ã states of C_{3}, C_{3}-Ne and C_{3}-Ar have been measured under supersonic molecular beam conditions. To minimize possible collisional quenching, the lifetimes of all three species were measured at a distance of about 42 nozzle diameters from the orifice. For all the vibrational levels of the Ã state of the C_{3} monomer, only J′=1 lifetimes were recorded, using R(0) lines. The accuracy of the lifetimes obtained was estimated to be better than 3ns. General features of the lifetimes of the C_{3} monomer are as follows: with increasing excitation of the symmetric stretching vibration, the lifetimes increase to a maximum at v_{1}=4. Bending excitation reduces the lifetimes, though exceptions occur for the 0 v 0 and 1 v 0 levels. Among all the vibrational levels, 0 2^{−} 0, 0 4^{−} 0, and 0 2^{+} 0 have the shortest lifetimes, similar to that of the origin level ^{a}. Lifetimes of the Ã state of C_{3}-Ne and C_{3}-Ar have only been measured for features 1.5-2 and 11-14 cm^{−1}, respectively, to the red of the R(0) lines of the Ã−X̃ bands of free C_{3}. It is not possible to reduce the backing pressure while measuring lifetimes of the complexes; we therefore reduced the average pressure of the chamber to 1-2x10^{−5} torr by lowering the repetition rate. No sign of predissociation was found in the Ne complex except for the level 0 12^{+} 0. The lifetimes of the complexes did not vary with vibration in the manner found for C_{3} itself. The lifetime of the 0 2^{−} 0 level, one of the lowest vibrational levels of the Ã state, was not affected by complexing with either Ne or Ar, consistent with the observation that the least van der Waals shifts were observed for this vibrational level ^{b}.

aG. Zhang, K.-S. Chen, A. J. Merer, Y.-C. Hsu, W.-J. Chen, S. Shaji, and Y.-A. Liao, J. Chem. Phys. 122, 244308 (2005).

bG. Zhang, B.-G. Lin, S.-M. Wen, and Y.-C. Hsu, J. Chem. Phys. 120, 3189 (2004).